OpenStudy (anonymous):

2x - 24 ________ = 2x -3 8

jimthompson5910 (jim_thompson5910):

what did you get for x

jimthompson5910 (jim_thompson5910):

If you're not sure how to start, you can multiply both sides by 8 to get 2x - 24 = 8(2x - 3)

OpenStudy (anonymous):

x - 24 over 4?

jimthompson5910 (jim_thompson5910):

2x - 24 = 8(2x - 3) 2x - 24 = 16x - 24 keep going

OpenStudy (anonymous):

x - 14 over 4

jimthompson5910 (jim_thompson5910):

no that's not correct

jimthompson5910 (jim_thompson5910):

2x - 24 = 16x - 24 2x-16x = -24+24 -14x = 0 x = 0/(-14) x = 0

OpenStudy (anonymous):

thats not an option

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what are your options and if possible, can you draw out the original problem? thanks

OpenStudy (anonymous):